OpenStudy (anonymous):

Factor the algebraic expression (x - 1)2 - (y - 2)2.

OpenStudy (anonymous):

\[a^2 - b^2 = (a+b)(a-b)\]

OpenStudy (anonymous):

\[(x+1)^2 - (y-2)^2 = [(x+1)+(y-2)] [(x+1)-(y-2)]\]